Operations Report – December 10th

Crew 184 Operations Report 10/12/2017

SOL: 9

Name of person filing report: HuntJ

Non-nominal systems: Generator non-functional, repaired and returned to operations in afternoon. (Operational Again!!!)

Notes on non-nominal systems: Mission Support found a solution with the generator today. It turns out that the ground and positive connectors to the battery were loose. Tightening them and a voltimeter check proved the battery was just fine. One more attempt at using jumper cables and the generator is back and running! Forward plan now is to leave it on until tomorrow morning (with fingers crossed that it stays on through the night). A decision will be made tomorrow morning on whether to turn the generator back off or leave it running.

Generator (hours run): Back on around 2:00pm. Running 3 hours and still going!

Solar— SOC @ 4:00pm: 99%

Diesel – 100%

Propane – 70 Percent Volume

Ethanol Free Gasoline (5 Gallon containers for ATV) – 1.5 Gallons

Water (trailer) – 600. gallons

Water (static) – 250 gallons

Trailer to Static Pump used – no

Water (loft) – Static to Loft Pump used – yes

Water Meter: 47 gallons

Toilet tank emptied: No

ATV’s Used: None,

Oil Added: No

ATV Fuel Used: None

# Hours the ATVs were Used today: 00:00 Hours

Notes on ATVs: No ATVs used today but were turned on for status checks. All nominal except for the Yamaha 300 which was determined a few days ago to have a dead battery.

Deimos rover used: Yes

Hours: 01:00 Hours

Beginning charge: 100%

Ending charge: 88.8%

Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR


Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: Yes

Hours: 01:00

Beginning charge: 100%

Ending charge: 51%

Currently charging: Yes

Opportunity rover used: Yes

Hours: 01:00 Hours

Beginning charge: 100%

Ending charge: 23%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 00:30 Hours

Beginning charge: 100%

Ending charge: 59%

Currently charging: Yes

HabCar used and why, where? NO

General notes and comments: Power is back and we’re back in sim!!

Summary of internet: 1.0 GB of Bandwidth

Summary of suits and radios: Radios are nominal after getting charged again.

EVA Suits:

EVA 6: Fan does not turn on.

EVA 8: Latches that secure the back panel are loose. Engaging them again does not change this. Concern is that back panel would fall off during EVA.

We will try to troubleshoot some of these issues and will update with changes.

Summary of Hab operations: Nominal

Summary of GreenHab operations: See Greenhab report

Summary of ScienceDome operations: Currently being cleaned

Summary of RAM operations: Not Operational

Summary of health and safety issues: All Healthy

Questions, concerns and requests to Mission Support:

Request more ATV Fuel. Can probably get through 2 more EVAs before exhausting remaining reserves.
Non burnable trash in airlock ready for pickup.

GreenHab Report – December 10th

GreenHab Report
Trisha Randazzo
December 10, 2017

Environmental control: (Choose with is appropriate and explain further if needed)

Ambient (no heating or/cooling)
high: 44
low: 21

Ambient with window/door open



Both heating and cooling

Shade cloth on/off: On

Average temperature: (N/A until new sensor is delivered) N/A

Changes to crops: Note all emergence of seedlings, death of seedlings, etc

New seedlings in the planter beds are doing well!
Yesterday five of the tomato plants that were thinned were doing badly, today two of them are in good health, two of them are on the mend, one is still doing poorly.

Daily water usage for crops: 2 gallons

Time(s) of watering for crops: 15:57

Morning research observations: (Currently not operational)

Changes to research plants: N/A

Daily watering and amount of water used:

Aquaponics: (Currently not operational) N/A

Narrative: Any other information you want to share

Continue to thin tomato plants
Finish planting herb seeds
Install CO detector and alarm
Changed automatic heater turn on temperature to 10 C
Plugged lights back in

Support/supplies needed:

Awaiting more soil and nutrients

Sol Summary – December 10th

MDRS Daily Summary Report for Sol 9

Summary Title:  Return of the Sim!

Mission Status:

Yesterday and today has been been primarily ‘out of sim’ due to the numerous HAB failures, including our evacuation last night due to a loss of power.  We spent the morning today working to repair the generator and appear to have gotten it in working order, though further work is necessary to ensure it’s long term successful operation (see Shannon’s email for details).  The solar array system is still misconfigured so power problems may yet pop up during the rest of our sim but we will take them as we go and troubleshoot, as a real martian crew would.

This afternoon we took the rovers out to drain the batteries per Mission Control request, once returned to the HAB we re-entered SIM at roughly 5pm.

It’s been a rough first week, but we are looking forward to getting back into sim and finishing out our mission successfully.

Sol Activity Summary:

1.      Out of SIM maintenance work

2.     General HAB cleaning / reports

Look Ahead Plan:

  1. Matryoshka EVA Site #5
  2. ‘Bricks’ communication experiment
  3. Solar Observation
  4. CPR Techniques
  5. Exercise
  6. General HAB Cleaning / Maintenance / Reports

Anomalies in work:

1.      HAB Leak from under EVA door continuing (per Mission Control leak is expected and has been present for months, presumed to be from Kitchen piping)

2.  Power System Misconfiguration ongoing

3.  2 EVA suit fans were non-functional during HAB check today, which may be related to power issues and resulting in them not being charged.  We have connected the suits to power and will assess tomorrow.  

Weather: Cold, but temps are above freezing in mid-day

Crew Physical Status: All crew in good physical health.


Reports to be filed:



EVA Request

Sol Summary

Support Requested:

1.      ATV Fuel

2.     Food Re-Supply

    1. 1 Box Raisin Bran
    2. 1 Box Powdered Milk
    3. 24 packs Instant Ramen
    4. 3 Boxes Pasta (any type)
    5. 2 Box Granola Bars
    6. 1 Box Saltine Cracker
    7. 1 Box Cracker (any type)
    8. 1 Box Hot Chocolate
    9. 1 Box Red Lobster Biscuits
    10. 1 Box Earl Grey Tea
    11. 1 Pack Chocolate Chips
    12. 1 Pack Tortillas
    13. 1 Pack Sour Cream\
    14. 1 Can SPAM
    15. 2 Can Chicken
Copyright © The Mars Society. All rights reserved. | Main Site